OpenStudy (anonymous):

how do I Simplify the following expression; express in terms of i. sqrt -54

OpenStudy (unklerhaukus):

\[\sqrt{-54}\] re write the \(-54\) as a product of factors \(-54=-1\times54=-1\times6\times9=-1 \times2\times3^3\) \[\sqrt{-54}\rightarrow\sqrt{-1 \times2\times3^3}\]

OpenStudy (unklerhaukus):

\[\sqrt{-1 \times2\times3^3}=\sqrt {-1}\times\sqrt 2\times\sqrt{3^3}\]\[=\sqrt {-1}\times\sqrt 2\times\sqrt{3\times3^2}\]\[=i\times\sqrt 2\times3\sqrt{3}\]\[=3\sqrt6i\]

OpenStudy (anonymous):

ok thank you I had up to the i x correct then i messed it up I see how you got it appreceiate it thank you